Sheridan Inc. has issued three types of debt on January 1, 2017, the start of the companys fiscal year (a) $12 million, 9-year, 14% unsecured bonds, interest payable quarterly. Bonds were priced to yield 12% (b) $26 million par of 9-year, zero-coupon bonds at a price to yield 12% per year (c) $20 million, 9-year, 11% mortgage bonds, interest payable annually to yield 12% Prepare a schedule that identifies the following items for each bond: (1) maturity value, (2) number of interest periods over life of bond, (3) stated rate per each interest period, (4) effective-interest rate per each interest period, (5) payment amount per period, and (6) present value of bonds at date of issue. (Round stated and effective rate per period to 2 decimal places, e.g. 10.25%. Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answer to 0 decimal places e.g, 58,971.) Unsecured Bonds Zero-Coupon Bonds Mortgage Bonds (1) Maturity value (2) Number of interest periods (3) Stated rate per period (4) Effective rate per period (5) Payment amount per period (6) Present value
Answer #1

  unsecured bond Zero coupon bond Mortgage bond

Maturity value $12million $26million $20million

Number of interest period 9*4=36    - 9

Stated rate per period 14/4=3.5% - 11%

Effective interest rate 3.14% 12/4=3% 12/4=3%

Payment amount per period   $420000 - $2200000

Present value $13309905 $9375860 $18934350

a) $3.309 1oS cosupon bend Yaues 6o00s0oX.3606 9 375 860 Reentra 1221SO+ 72122 터.fective Rate ) Dnse cused bond no oㄔ cchive hat pe peod12.SS314 1 Wate value fachous cam be caledateod wFh